OpenStudy (anonymous):

At the Olympics, the U.S. won 14 more gold than silver medals in Sydney, Australia in 2000. They won a total of 64 gold and silver medals. How many gold medals did the U.S. win?

OpenStudy (javk):

wrong 64 is the total

OpenStudy (anonymous):

oh sorrry

rishavraj (rishavraj):

let total no of silver medals won = x so no of gold medals won = 14 + x it means x + (14 + x) = 64

OpenStudy (anonymous):

x=25 for your problem but thats not one of my choices

rishavraj (rishavraj):

u have been asked about no of gold medals...not the silver one... x = no of silver medal x + 14 = no of gold medal`

rishavraj (rishavraj):

@mathsciencehistory

OpenStudy (anonymous):

At the Olympics, the U.S. won 14 more gold than silver medals in Sydney, Australia in 2000. They won a total of 64 gold and silver medals. How many gold medals did the U.S. win? 36 37 39 41

rishavraj (rishavraj):

whts 25 + 14 ????? @mathsciencehistory

OpenStudy (anonymous):

oh lmao